OpenStudy (anonymous):

Which of the following is a polynomial with roots -√(3), √(3), and 2? x^3 – 2x^2 – 3x + 6 x^3 + 2x^2 – 3x – 6 x^3 – 3x^2 – 5x + 15 x^3 + 3x^2 – 5x – 15

OpenStudy (anonymous):

Expand y = (x - x1)(x - x2)(x - x3) where x1, x2, and x3 are your roots. Make this substitution and multiply out.

OpenStudy (anonymous):

can u show me how to do that?

OpenStudy (anonymous):

y = (x - x1)(x - x2)(x - x3) = [x + sqrt(3)][x - sqrt(3)](x - 2) and then expand that.